# Quadratic counts of highly tangent lines to hypersurfaces

### v1

@@ 5. Highly tangent lines and the second fundamental form
(see e.g. [KW21, p. 692]). By base changing to $k(p)$, working $k(p)$-rationally, and then applying the field trace, we may assume (without loss of generality) that all zeros of   II are $k$-rational.
Equation (5.5) only makes sense once we have chosen a Nisnevich coordinate on $X$ and a compatible local trivialization of $\mathcal{F}$. We will explain why our prior argument is independent of our choice of coordinate and trivialization. By [KW21, Proposition 20], we always have Nisnevich coordinates $(U,\varphi)$ about $p$, where $\varphi\colon U\to\mathbb{A}^{1}$. By intersecting with our affine neighborhood $\operatorname{Spec}(A)$, we may assume that $U=\operatorname{Spec}(A)$. The map $\varphi$ corresponds to a local parameter $t$ of $\operatorname{Spec}(A)$, and $\mathrm{d}{t}:=t\otimes 1-1\otimes t$ will generate the ideal $I$. We may therefore take $z$ to be the local parameter determined by our Nisnevich coordinate. Finally, one can always obtain a trivialization of $\mathcal{F}$ compatible with $(U,\varphi)$ by [McK22, Proposition 5.5]. In the present case, this can be done by picking a non-vanishing local section of the line bundle $\mathcal{F}$ and rescaling if necessary. ∎
@@
Higher fundamental forms for algebraic varieties made their first appearance in work of Griffiths and Harris [GH79], and were later generalized by several authors [Teg92, Lan98, EN23, MP24]. It would be interesting to try extending the results of this section to hypersurfaces in higher-dimensional projective spaces. One of the first hurdles to such a generalization is that the sheaves used to define higher fundamental forms need not be vector bundles.

### v2

@@ 5. Highly tangent lines and the second fundamental form
(see e.g. [EnrichedCubicSurface, p. 692]). By base changing to $k(p)$, working $k(p)$-rationally, and then applying the field trace, we may assume (without loss of generality) that all zeros of   II are $k$-rational.
Equation (5.5) only makes sense once we have chosen a Nisnevich coordinate on $X$ and a compatible local trivialization of $\mathcal{F}$. We will explain why our prior argument is independent of our choice of coordinate and trivialization. By [EnrichedCubicSurface, Proposition 20], we always have Nisnevich coordinates $(U,\varphi)$ about $p$, where $\varphi\colon U\to\mathbb{A}^{1}$. By intersecting with our affine neighborhood $\operatorname{Spec}(A)$, we may assume that $U=\operatorname{Spec}(A)$. The map $\varphi$ corresponds to a local parameter $t$ of $\operatorname{Spec}(A)$, and $\mathrm{d}{t}:=t\otimes 1-1\otimes t$ will generate the ideal $I$. We may therefore take $z$ to be the local parameter determined by our Nisnevich coordinate. Finally, one can always obtain a trivialization of $\mathcal{F}$ compatible with $(U,\varphi)$ by [McK22, Proposition 5.5]. In the present case, this can be done by picking a non-vanishing local section of the line bundle $\mathcal{F}$ and rescaling if necessary. ∎
@@
Higher fundamental forms for algebraic varieties made their first appearance in work of Griffiths and Harris [MR559347], and were later generalized by several authors [Teg92, landsbergseoul, einniu, MP24]. It would be interesting to try extending the results of this section to hypersurfaces in higher-dimensional projective spaces. One of the first hurdles to such a generalization is that the sheaves used to define higher fundamental forms need not be vector bundles.